The sum of the ages of Toni and Jaime is 60. Nine years ago, Toni was twice as old as Jaime. How old are they now?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39620)
Answer by ikleyn(52799)

.

Nine years ago, Toni was twice as old as Jaime, AND, at the same time,  the sum of their ages was 60-2*9 = 42.


Hence, 9 years ago they were 14 (Jaime) and 28 (Toni).


Hence, now they are 14+9 = 23 (Jaime) and 28+9 = 37 (Toni) years old.    ANSWER